A triangular field is surrounded by 470 yards of fence. The longest side of the feild is 10 yards less than three times the length of the shortest side. The length of the next longest side is two times the length of the shortest side. What is the length of the shortest side?
Answer by oberobic(2304)

A + B + C = 470 yd
A = shortest side
C = longest side
C = 3A -10
B = 2A
.
substitute
.
A +2A +3A-10 = 470
.
6A = 480
.
A = 80
.
B = 2A = 160
.
C = 3A-10 = 240 - 10 = 230
.
P = 80+160+230 = 470
Correct
.
Answer: The shortest side =is 80 yd.
.
Done.